OpenStudy (anonymous):

make a conjecture about the next item in the sequence 3,6,12,24... 32 36 48 72 answer pleaseee!!!

OpenStudy (anonymous):

3*2=6*2=12*2=24*2=...

OpenStudy (primeralph):

Well, 6 is double of 3 and 12 is double of 6, and 24 is double of 12, so what is double of 24?

OpenStudy (anonymous):

24 multiplied by two would be what?

OpenStudy (anonymous):

48

OpenStudy (primeralph):

Yeah, there you go.

OpenStudy (anonymous):

Yes, the answer would be 48